OpenStudy (kitkat1):

What is the volume of a sphere with a radius of 9 inches? 108 242 642 972

OpenStudy (anonymous):

3053.63in³

OpenStudy (kitkat1):

There is no such a answer

OpenStudy (anonymous):

3053.63in^3

OpenStudy (kitkat1):

look at the answers

OpenStudy (kitkat1):

which one from the chart

OpenStudy (anonymous):

Radius is half... Give me a sec I am writing it down.

OpenStudy (anonymous):

@hislightningbug812 is correct, but the answer is not up there.... I don't get it... The answer is infact, V≈3053.63

OpenStudy (anonymous):

Sorry V=3053.63

OpenStudy (kitkat1):

there is no such a answer in my answers.

OpenStudy (anonymous):

Well that is the answer.. Sorry dude, but that is it.

OpenStudy (anonymous):

Two people got the same answer.

OpenStudy (anonymous):

If it helps, it is 972 cubic meters.

OpenStudy (anonymous):

So maybe your teacher typed it wrong and it is supposed to be in meters?

OpenStudy (kitkat1):

the answers are like this 108 pi 242 pi 642 pi 972 pi

OpenStudy (anonymous):

The answer is 972 pi

OpenStudy (anonymous):

you didn't put the pi's in Had you I would have gave you the answer right away.

OpenStudy (triciaal):

volume of a sphere V = 4/3*pi*r^3 using pi as 22/7 and given r = 9 V = 4/3*22/7*9^3 V =

OpenStudy (anonymous):

@CMFink00 I did that so I can come back and look at comments in a minute.

OpenStudy (kitkat1):

okay thanks for that can u help me with one more?

OpenStudy (anonymous):

Sure post it.

OpenStudy (kitkat1):

What is the length of side AB of the triangle?

OpenStudy (triciaal):

yes 972 pi

OpenStudy (anonymous):

What other information do you have? Like numbers or something? I cannot answer it without them.

OpenStudy (kitkat1):

Answers are 3, 6, 6 in exponent and 18 in exponent.

OpenStudy (anonymous):

I am pretty sure your answer is 3.

OpenStudy (triciaal):

AB is the hypotenuse of a right triangle AB^2 = AC^2 + BC^2

OpenStudy (triciaal):

do you know how to calculate the lengths of each side AC and BC given the coordinates of A, B and C?

OpenStudy (anonymous):

I have to go do my homework Bye.

OpenStudy (kitkat1):

Does anyone else agree with 3?

OpenStudy (anonymous):

Yes. 3 is the answer.

OpenStudy (anonymous):

Count the points going up the side of ab. there are 3.

OpenStudy (kitkat1):

I need someone to confirm

OpenStudy (triciaal):

disagree strongly

OpenStudy (kitkat1):

what do u think triciaal?

OpenStudy (anonymous):

It's like counting wavelengths... I don't know if you know how to do that. I am a freshman in highschool. I don't know what grade you are in. Sounds like 7th.

OpenStudy (kitkat1):

im in 8th

OpenStudy (triciaal):

@Kitkat1 I think you are not reading my post

OpenStudy (kitkat1):

which one?

OpenStudy (triciaal):

repost 1) AB is the hypotenuse of a right triangle AB^2 = AC^2 + BC^2 repost 2) do you know how to calculate the lengths of each side AC and BC given the coordinates of A, B and C? take a pick

OpenStudy (kitkat1):

So u dont agree with the 3 answer?

OpenStudy (kitkat1):

I think I can use the pythagorean theorem for the second one right?>

OpenStudy (triciaal):

repost 3) no

OpenStudy (kitkat1):

then what is it

OpenStudy (triciaal):

repost 1) AB is the hypotenuse of a right triangle AB^2 = AC^2 + BC^2 this is pythagorean theorem

OpenStudy (kitkat1):

Yes but im already sitting on this quiz for about an hour and can't find the answer for it what is it?

OpenStudy (triciaal):

are you mixing up the radicand sign with exponent ? an exponent is what you raise a base to